OpenStudy (xhaleymariexo):

I really need help with this, I don't want the answer I really need help solving this. The expression 1000(1.1)^t represents the value of a $1000 investment that earns 10% interest per year, compounded annually for t years. What is the value of a $1000 investment at the end of 3 years? So far I have 1000(1.1)^3, I have no idea where to go from here.

OpenStudy (will.h):

Interest exponential function is F(x) = p(1 + r)^x We have p = 1000 R= 10/100 because it is 10% And x represents the number of years which is 3 so therefore apply the right values and.youll get your answer :)

jhonyy9 (jhonyy9):

(1,1)^3 = 1,1*1,1*1,1 = 1,331 1000*1,331 = 1331 hope helped

OpenStudy (will.h):

Let me know if you have further questions

OpenStudy (will.h):

And yeah the answer is 1331

OpenStudy (phi):

I would use a calculator 1.1 "cubed" or 1.1 to the 3 power i.e. 1.1^3 or 1.1*1.1*1.1 then multiply by 1000 or type into google 1000(1.1)^3=